In a scale drawing of a house, the living room is 1 in. long. The actual lenght is 24 ft.
The actual width of the living room is 18 ft. Find the width on the drawing.

Answer by checkley79(3341)

1/24=X/18 CROSS MULTIPLY.
24X=18
X=18/24
X=3/4 OR .75 INCHES.